OpenStudy (anonymous):

integrate: 0 to infinity ʃ e^tx 1/6(e^-x/2 +xe^-x/2)

OpenStudy (zarkon):

Finding a Moment Generating Function? distribute and integrate term by term.

OpenStudy (zarkon):

I get \[\frac{3-2t}{3(2t-1)^2}\] for \[t<1/2\]

OpenStudy (anonymous):

yes mgf

OpenStudy (anonymous):

how did u know?

OpenStudy (zarkon):

I've done a lot of these...could tell by just looking at it.

OpenStudy (anonymous):

why is it t<1/2 i dont undertsand

OpenStudy (zarkon):

\[\int\limits_{0}^{\infty}e^{tx}e^{-x/2}dx\] \[\int\limits_{0}^{\infty}e^{tx-x/2}dx\] \[\int\limits_{0}^{\infty}e^{(t-1/2)x}dx\] this integral will only converge when (t-1/2)<0 ie t<1/2

OpenStudy (anonymous):

so did u take 1/6 out first, then expamded the bracckets and multiplied e^tx with the two exponetials insdie the brackets. is it possible to show me the steps pleeez? :)

OpenStudy (zarkon):

yes...pull the 1/6 outside the integral...then distribute e^{tx}...then integrate it term by term...you can do it.

OpenStudy (anonymous):

ok i will post my answer..thanks

OpenStudy (anonymous):

do u have ti integrate xe^(t-1/2)x by parts?

OpenStudy (zarkon):

yes

OpenStudy (anonymous):

thnx

OpenStudy (zarkon):

no problem